2. Calculate the pH of a 0.36 M CH,COONa solution. The K = 5.6X 10-10
Answer #1

The solution of such salt is basic in nature . The anion of the salt is reactive . It racts with water to form a weak acid and OH- .

CH3Co0 — (аg) + H2O = CH3COOH + OH-

pH = 7 + 1/2 pKa + 1/2 logC

Also, pKa + pKb = 14

Kb = 5.6*10^-10 , pKb = - log Kb = - log ( 5.6 * 10^-10) = 9.25

pKa = 14- 9.25 = 4.748

Therefore, pH = 7+ 4.748/2 + log(.36)/2

pH = 7+ 2.374- .221= 9.153